OpenStudy (anonymous):

What is the percent of increase from 60 to 75? A. 80% B. 25% C. 20% D. 15%

OpenStudy (anonymous):

I think its 25 ;-; Im not sure tho.

OpenStudy (kendricklamar2014):

60 + 15% = 69 60 + 20% = 72 60 + 25% = 75

OpenStudy (anonymous):

the answer is b

OpenStudy (kendricklamar2014):

Don't give Direct Answers @teetee4455

OpenStudy (anonymous):

Thx @KendrickLamar2014

OpenStudy (kendricklamar2014):

\[No~Problem\]

OpenStudy (freckles):

first what is the change from 60 to 75 then once you find the change divide the change by what you started with

OpenStudy (anonymous):

ohh my bad @KendrickLamar2014

OpenStudy (freckles):

\[\frac{x_2-x_1}{x_1} \text{ will give you your percent increase }\]

OpenStudy (freckles):

\[\text{ where } x_2 \text{ is your number after the change happened } \\ \text{ and } x_1 \text{ was that initial number }\]

OpenStudy (anonymous):

Thanks for explaining @freckles

OpenStudy (freckles):

np